# Bigram Subnetworks: Mapping to Next Tokens in Transformer Language Models

### v2

@@ 1 Introduction
[…]
@@ 2 Background and Related Work
There is good reason to expect Transformer language models to exhibit bigram subnetworks, based on previous research drawing connections between Transformer language models and traditional $n$-gram models. An $n$-gram model predicts each next token $w_{i}$ based only on the previous $n-1$ tokens: $P(w_{i}|w_{i-1},...w_{i-n+1})$. Transformer models trained on formal languages or directly on $n$-grams can learn $n$-gram distributions (Elhage et al., 2021; Svete & Cotterell, 2024; Svete et al., 2024), and they can recognize $n$-grams in context using specialized attention heads (Akyürek et al., 2024). When trained on natural language, Transformer models overfit to $n$-gram probabilities for increasing $n$ early in pretraining (Chang & Bergen, 2022; Choshen et al., 2022; Chang et al., 2024), and their predictions can be approximated to some degree using $n$-gram rule sets (Nguyen, 2024). Smoothed $n$-grams alone achieve fairly high next token prediction accuracy for large $n$ (Liu et al., 2024).
@@ 3.1 Method: Continuous Sparsification
$\textrm{Loss}(M,x)=\textrm{CrossEntropy}\Big{(}P(x),\textrm{MaskedModel}_{M}(x)\Big{)}+\lambda\frac{||M||_{1}}{|M|}$ (1)
As in Savarese et al. (2020), $\lambda$ weights the L1 penalty term $||M||_{1}$, which enforces sparsity in the mask $M$. We normalize by the total number of trainable mask parameters $|M|$, and we train subnetworks for $\lambda\in[0,1,5,10,50,100,500,1000]$ to evaluate the effects of sparsity on subnetwork performance. We learn the mask $M$ over all model parameters except input and output embedding parameters and layernorm parameters. We train each subnetwork on sequences of 128 tokens with batch size 32 and learning rate 5e-5, and we set the mask sigmoid temperature to divide by 1.001 per training step (dictating how fast $M$ approaches a binary mask). We find that these hyperparameters generally allow subnetwork training to converge for all models tested, and the resulting subnetworks have similar evaluation loss to subnetworks trained under different hyperparameter settings. Subnetwork training and convergence details are in §A.2.
@@ 3.2 Existence of Bigram Subnetworks
As shown in Figure 1 (right), for all models, continuous sparsification is able to find subnetworks whose surprisals (i.e. token-level losses or negative log-probabilities) correlate highly with those of bigrams (Pearson’s $r>0.95$). We measure surprisal correlations rather than cross-entropies themselves because they are more interpretable across models, they are more efficient to compute when bigram surprisals can be cached, and we find that they produce similar patterns of results. Notably, we find that bigram correlations plateau at roughly 10M active non-embedding parameters regardless of model size (Figure 1, left, center). [^4] In §A.3, we discuss why we would expect bigram correlations to plateau with more allowed parameters. Of course, as subnetworks approach the full models, bigram correlations will return to the lower full model correlations (Figure 1, right). This suggests that bigram subnetwork “size” is roughly independent of full model size, and thus the bigram subnetwork comprises a much smaller proportion of model parameters in larger models. For example, in Pythia 1B, a subnetwork containing only 0.17% of non-embedding parameters reaches a bigram surprisal correlation of $r=0.959$.
@@ 4.1 Bigram Subnetworks Persist but Decompress Throughout Pretraining
Figure 2: Left: estimated bigram surprisal correlation for bigram subnetworks with different numbers of active parameters (excluding embedding parameters) for Pythia 1B at different checkpoints (§4.1). Center, right: proportions of parameters in the Pythia 1B bigram subnetwork that are in each MLP and attention layer throughout pretraining (§4.2). Note that the color bar scale is 10$\times$ larger for MLP proportions, as a far greater proportion of bigram subnetwork parameters are in the MLP layers.
@@ 5.1 Rotations Induced by the First Transformer Layer
For each layer, we compute activation vectors $X_{\ell}\in\mathbb{R}^{n\times d}$ outputted by that layer for $n=$128K tokens in context, using sequences from the OSCAR dataset as in §3. For each layer, we use ridge regression to find a linear map $L_{in}\in\mathbb{R}^{d\times d}$ that maps activations in layer $\ell$ to their original input token embeddings $X_{in}$, and a linear map $L_{out}\in\mathbb{R}^{d\times d}$ that maps activations in layer $\ell$ to their corresponding output layer activations $X_{out}$. [^7] We use ridge regression because the correlations between different activation dimensions make linear regression coefficients (i.e. entries of $L$) unstable without normalization. We use L2 regularization term $\alpha=1$. We then assess the structure of these linear transformations. A linear transformation $L\in\mathbb{R}^{d\times d}$ has $d$ complex eigenvalues including multiplicity. Because $L$ has only real entries, the eigenvalues appear in complex conjugate pairs $a\pm bi$. Each eigenvalue magnitude $||a\pm bi||$ corresponds to scaling a given input dimension, and the corresponding eigenvalue angle ($\textrm{arctan}(b/a)$) corresponds to rotating that input dimension. We focus on the rotations induced by each linear transformation $L$, because scalings are intuitively less consequential for representing features in activation space; if a given feature is represented along a direction $v$, then scaling that dimension will still encode the same information along the same direction, just multiplied by a scalar. Thus for each layer in each model, we compute the median rotation $r\in[0,180]$ degrees for the transformation $L_{in}$ to input activations and for the transformation $L_{out}$ to output activations. Loosely, these median rotations quantify the “transformation distance” from a layer’s activations to current vs. next token representations.
@@ 6.2 Optimal Subnetworks Overlap Significantly with Bigram Subnetworks
We report results for Pythia 1B here, and we report similar results for other models in §A.9. In Pythia 1B, the bigram subnetwork and optimal subnetwork parameter overlap is 15.3$\times$ greater than would be expected from chance ($p<0.0001$). In that model, 38.0% of the bigram subnetwork is contained in the optimal subnetwork, even though the optimal subnetwork is only 0.93% of model parameters and the bigram subnetwork is 0.17% of model parameters (excluding embedding parameters). This indicates that sparse optimal subnetworks are not only similar to bigram subnetworks in behavior, but they also identify similar individual model parameters.
@@ 7 Discussion and Conclusion
These properties also make bigram subnetworks ideal starting points from which to build language models up from component circuits. By adding specific circuits to bigram subnetworks and evaluating how model predictions change from bigram predictions, researchers can evaluate the sufficiency of those circuits for performing hypothesized functions. Notably, due to the sparsity of the bigram subnetworks and the relative simplicity of the bigram function, we can be more confident that the added circuits are leveraging their own suboperations and intermediate features rather than scaffolding off of preprocessed features in an opaque language model. To facilitate these lines of research, we publicly release bigram subnetworks for GPT-2 and Pythia models up to 1B parameters. As researchers un-ablate interpretable circuits, we can aim to shrink the performance gap between fully interpretable model subnetworks (i.e. subnetworks built up from interpretable circuits) and their corresponding full models, setting concrete targets for mechanistic interpretability research.
@@ A.1 Circuit Necessity, Sufficiency, Faithfulness, Completeness, and Minimality
  • For a circuit $C\subseteq$ model $M$, circuit necessity requires that ablating $C$ from $M$ removes the target behavior from $M$. Ideally, if $M$ has some more complex behavior $f_{0}$, and $C$ is hypothesized to perform target behavior $f_{1}$, then $M\backslash C$ should ideally perform $f_{0}-f_{1}$ if there is some intuitive and well-defined way to “subtract” behavior $f_{1}$ from $f_{0}$.
@@ A.2 Subnetwork Training Details
As described in §3.1, we use continuous sparsification (Savarese et al., 2020; Lepori et al., 2023b) to identify subsets of model parameters that minimize cross-entropy loss with the bigram distribution. To do this, we build upon the implementation in Lepori et al. (2023a). For all fully trained models (§3) and when training “optimal” subnetworks (§6), we train subnetworks with sparsity terms $\lambda\in[0,1,5,10,50,100,500,1000]$. For checkpoints during pretraining (§4), we train subnetworks with $\lambda\in[0,1,10,100,500]$. Subnetworks are trained on sequences of 128 tokens with batch size 32 and learning rate 5e-5, with a multiplicative sigmoid temperature decrease of 1.001 per training step (dictating how fast the continuous mask approaches a binary mask). We learn the subnetwork mask over all model parameters except the token embedding, token unembedding, and layernorm parameters. All mask parameters are initialized at $0.0$ before the sigmoid ($0.50$ after the sigmoid).
For each subnetwork, we stop training when the number of mask parameters in $[0.10,0.90]$ (i.e. “undecided” mask parameters) is less than 1% of the number of mask parameters greater than $0.90$ (i.e. parameters to be included in the subnetwork). This ensures that the number of “undecided” parameters will not change the learned subnetwork by more than 1% when the continuous mask is converted to a binary mask. In practice, it takes roughly 5K to 10K training steps to reach this convergence criterion. To verify training stability, we check for any spikes in cross-entropy loss of $+0.25$ or ×$\times 1.25$. If such a spike is not recovered at least 100 steps before the end of training, we re-train the subnetwork. Each subnetwork takes approximately four to twelve hours to train on a single NVIDIA RTX A6000 (48GB) GPU, depending primarily on model size. Subnetwork training costs make up the vast majority of computational costs to produce the results in this paper.
@@ A.3 Bigram Subnetwork Selection
Thus in our analyses in §4.2, §5, and §6, for GPT-2 (re-trained), GPT-2 small, and Pythia 160M, we use the bigram subnetwork trained with $\lambda=10$. For GPT-2 large and Pythia 1B, we use the bigram subnetwork trained with $\lambda=100$. The parameter counts and bigram surprisal correlations for these subnetworks are reported in Table 1. As noted in §6, for optimal subnetworks, we consider the sparsest optimal subnetwork that still contains more active parameters than the bigram subnetwork selected for that model. In general, we find that our results remain qualitatively similar for different $\lambda$, as long as the bigram subnetworks are reasonably sparse (e.g. $<10\%$ of parameters active).
@@
A.4 Importance of Trained Token Embeddings
However, keeping only the token embedding and unembedding parameters (and layernorm parameters) does not result in high bigram surprisal correlations (Table 2; “embeddings, empty subnetwork”). In other words, the embeddings do not recreate the $W_{U}W_{E}$ embedding-unembedding bigram structure described for zero-layer Transformers in Elhage et al. (2021). This is not entirely surprising, but it indicates that the bigram subnetwork parameters we find in Transformer layers are still necessary for high bigram correlations.
@@ A.6 Checkpoint Results for Other Models
Figure 6: Results as in Figure 2, but for GPT-2 small* and Pythia 160M. GPT-2 small* indicates the GPT-2 small replication from Chang et al. (2024) with checkpoints. Left: estimated bigram surprisal correlation for bigram subnetworks with different numbers of active parameters (excluding embedding parameters) at different checkpoints. Right: proportions of parameters in the subnetwork that are in each MLP and attention layer throughout pretraining. Note that the color bar scale is 5$\times$ or 10$\times$ larger for MLP proportions, as a far greater proportion of bigram subnetwork parameters are in the MLP layers.
@@ A.9 Optimal Subnetwork Results for Other Models
In Table 3, we report the actual and expected (by chance) parameter overlaps between the bigram subnetwork and the optimal subnetwork for each model, as described in §6.2 (with bigram subnetworks selected as in §A.3). As in Pythia 1B in the main text (§6.2), for all models, the actual overlaps between bigram subnetworks and optimal subnetworks are far larger than would be expected by chance.
@@
Pythia 160M 8.52e-2 1.33e-1 5.23e-2 1.37e-2 3.81$\times$ 61.4%
Pythia 1B 1.71e-3 9.30e-3 6.51e-4 4.26e-5 15.27$\times$ 38.0%
GPT-2 small* 1.09e-2 1.77e-2 2.41e-3 3.43e-4 7.03$\times$ 22.1%
GPT-2 small 1.07e-2 1.37e-2 1.53e-3 2.10e-4 7.27$\times$ 14.3%
GPT-2 large 1.03e-3 7.06e-3 3.34e-4 2.63e-5 12.70$\times$ 32.5%

### v3

@@ 1 Introduction
    Consist of roughly 10M parameters regardless of model size, consistently reaching bigram surprisal correlations of r>0.95r>0.953).
@@ 2 Background and Related Work
There is good reason to expect Transformer language models to exhibit bigram subnetworks, based on previous research drawing connections between Transformer language models and traditional $n$-gram models. An $n$-gram model predicts each next token $w_{i}$ based only on the previous n1n-1 tokens: $P(w_{i}|w_{i-1},...w_{i-n+1})$. Transformer models trained on formal languages or directly on $n$-grams can learn $n$-gram distributions (Elhage et al., 2021; Svete & Cotterell, 2024; Svete et al., 2024), and they can recognize $n$-grams in context using specialized attention heads (Akyürek et al., 2024). When trained on natural language, Transformer models overfit to $n$-gram probabilities for increasing $n$ early in pretraining (Chang & Bergen, 2022; Choshen et al., 2022; Chang et al., 2024), and their predictions can be approximated to some degree using $n$-gram rule sets (Nguyen, 2024). Smoothed $n$-grams alone achieve fairly high next token prediction accuracy for large $n$ (Liu et al., 2024).
@@ 3.1 Method: Continuous Sparsification
[…]
As in Savarese et al. (2020), $\lambda$ weights the L1 penalty term $M||_{1}$, which enforces sparsity in the mask $M$. We normalize by the total number of trainable mask parameters $|M|$, and we train subnetworks for $\lambda\in[0,1,5,10,50,100,500,1000]$ to evaluate the effects of sparsity on subnetwork performance. We learn the mask $M$ over all model parameters except input and output embedding parameters and layernorm parameters. We train each subnetwork on sequences of 128 tokens with batch size 32 and learning rate 5e-5, and we set the mask sigmoid temperature to divide by 1.001 per training step (dictating how fast $M$ approaches a binary mask). We find that these hyperparameters generally allow subnetwork training to converge for all models tested, and the resulting subnetworks have similar evaluation loss to subnetworks trained under different hyperparameter settings. Subnetwork training and convergence details are in §A.2.
@@ 3.2 Existence of Bigram Subnetworks
As shown in Figure 1, for all models, continuous sparsification is able to find subnetworks whose surprisals (i.e. token-level losses or negative log-probabilities) correlate highly with those of bigrams (Pearson’s $r>0.95$). We measure surprisal correlations rather than cross-entropies themselves because they are more interpretable across models, they are more efficient to compute when bigram surprisals can be cached, and we find that they produce similar patterns of results. Notably, we find that bigram correlations plateau at roughly 10M active non-embedding parameters regardless of model size (Figure 1, left, center). [^4] In §A.3, we discuss why we would expect bigram correlations to plateau with more allowed parameters. Of course, as subnetworks approach the full models, bigram correlations will return to the lower full model correlations (Figure 1, right). This suggests that bigram subnetwork “size” is roughly independent of full model size, and thus the bigram subnetwork comprises a much smaller proportion of model parameters in larger models. For example, in Pythia 1B, a subnetwork containing only 0.17% of non-embedding parameters reaches a bigram surprisal correlation of $r=0.959$.
@@ 4.1 Bigram Subnetworks Persist but Decompress Throughout Pretraining
[…]
@@ 5.1 Rotations Induced by the First Transformer Layer
For each layer, we compute activation vectors $X_{\ell}\in\mathbb{R}^{n\times d}$ outputted by that layer for n=n=128K tokens in context, using sequences from the OSCAR dataset as in §3. For each layer, we use ridge regression to find a linear map $L_{in}\in\mathbb{R}^{d\times d}$ that maps activations in layer $\ell$ to their original input token embeddings $X_{in}$, and a linear map $L_{out}\in\mathbb{R}^{d\times d}$ that maps activations in layer $\ell$ to their corresponding output layer activations $X_{out}$. [^7] We use ridge regression because the correlations between different activation dimensions make linear regression coefficients (i.e. entries of $L$) unstable without normalization. We use L2 regularization term $\alpha=1$. We then assess the structure of these linear transformations. A linear transformation $L\in\mathbb{R}^{d\times d}$ has $d$ complex eigenvalues including multiplicity. Because $L$ has only real entries, the eigenvalues appear in complex conjugate pairs $a\pm bi$. Each eigenvalue magnitude $a\pm bi||$ corresponds to scaling a given input dimension, and the corresponding eigenvalue angle ($\textrm{arctan}(b/a)$) corresponds to rotating that input dimension. We focus on the rotations induced by each linear transformation $L$, because scalings are intuitively less consequential for representing features in activation space; if a given feature is represented along a direction $v$, then scaling that dimension will still encode the same information along the same direction, just multiplied by a scalar. Thus for each layer in each model, we compute the median rotation $r\in[0,180]$ degrees for the transformation $L_{in}$ to input activations and for the transformation $L_{out}$ to output activations. Loosely, these median rotations quantify the “transformation distance” from a layer’s activations to current vs. next token representations.
@@ 6.2 Optimal Subnetworks Overlap Significantly with Bigram Subnetworks
[…]
@@ 7 Discussion and Conclusion
Speculatively, we hypothesize that bigram subnetworks emerge because early in training (or for low-performing models), bigram heuristics near-optimally decrease the language modeling loss overall, so the optimizer pushes parameters towards recreating a similar distribution. Thus, early in training (e.g. when the model overall approximates the bigram distribution), the primary function of the bigram subnetwork is likely to actually make a prediction similar to a bigram prediction. Later in training, we speculate that the model scaffolds off of these bigram predictions and subnetworks, adapting the early layers to perform generally useful early-layer computations (e.g. transforming to next token space) that still maintain some similarity to bigram predictions. It remains ambiguous whether these early-layer computations constitute “intermediate bigram predictions” (e.g. preliminary predictions before refinement by later layers; §A.5), or whether the bigram subnetwork becomes a more deeply integrated part of model processing, not separable as an “intermediate prediction”. Regardless, the persisting similarity to bigram-like computations in early layers is likely what makes it possible to extract bigram subnetworks even late in training, when these subnetworks have adapted to scaffold other computational roles as well.
Notably, even in fully trained models, the minimal and interpretable properties of bigram subnetworks make them ideal starting points from which to build language models up from component circuits. By adding specific circuits to bigram subnetworks and evaluating how model predictions change from bigram predictions, researchers can evaluate the sufficiency of those circuits for performing hypothesized functions. Due to the sparsity of the bigram subnetworks and the relative simplicity of the bigram function, we can be more confident that the added circuits are leveraging their own suboperations and intermediate features rather than scaffolding off of preprocessed features in an opaque language model. To facilitate these lines of research, we publicly release bigram subnetworks for GPT-2 and Pythia models up to 1B parameters. As researchers un-ablate interpretable circuits, we can aim to shrink the performance gap between fully interpretable model subnetworks (i.e. subnetworks built up from interpretable circuits) and their corresponding full models, setting concrete targets for mechanistic interpretability research.
@@ A.1 Circuit Necessity, Sufficiency, Faithfulness, Completeness, and Minimality
[…]
@@ A.2 Subnetwork Training Details
As described in §3.1, we use continuous sparsification (Savarese et al., 2020; Lepori et al., 2023b) to identify subsets of model parameters that minimize cross-entropy loss with the bigram distribution. We estimate the bigram distribution by counting bigram frequencies in 1.28B tokens of OSCAR web text (10M sequences of 128 tokens; Abadji et al., 2021), tokenized with the corresponding model tokenizer. Because the corpus is large enough that all tokens appear at least once (i.e. all possible bigram “prefixes”), there are no undefined bigram probabilities in our experiments, and we do not need to smooth the bigram distribution estimate.
To train the subnetworks themselves, we build upon the implementation in Lepori et al. (2023a). For all fully trained models (§3) and when training “optimal” subnetworks (§6), we train subnetworks with sparsity terms $\lambda\in[0,1,5,10,50,100,500,1000]$ (Equation 1). For checkpoints during pretraining (§4), we train subnetworks with $\lambda\in[0,1,10,100,500]$. Subnetworks are trained on sequences of 128 tokens with batch size 32 and learning rate 5e-5, with a multiplicative sigmoid temperature decrease of 1.001 per training step (dictating how fast the continuous mask approaches a binary mask). We learn the subnetwork mask over all model parameters except the token embedding, token unembedding, and layernorm parameters. [^10] We do not include embedding and unembedding (language model head) parameters in the mask because we hypothesize that embedding parameters for rare tokens might be highly likely to get masked. Specifically, when optimizing for bigram cross-entropy loss and encouraging sparsity, the continuous sparsification method might be likely to find it optimal to simply drop many embedding and unembedding parameters for rarer tokens, because they contribute less to the loss. All mask parameters are initialized at $0.0$ before the sigmoid ($0.50$ after the sigmoid).
For each subnetwork, we stop training when the number of mask parameters in $[0.10,0.90]$ (i.e. “undecided” mask parameters) is less than 1% of the number of mask parameters greater than $0.90$ (i.e. parameters to be included in the subnetwork). This ensures that the number of “undecided” parameters will not change the learned subnetwork by more than 1% when the continuous mask is converted to a binary mask. In practice, it takes roughly 5K to 10K training steps to reach this convergence criterion. To verify training stability, we check for any spikes in cross-entropy loss of +0.25+0.25 or $\times 1.25$. If such a spike is not recovered at least 100 steps before the end of training, we re-train the subnetwork. Each subnetwork takes approximately four to twelve hours to train on a single NVIDIA RTX A6000 (48GB) GPU, depending primarily on model size. Subnetwork training costs make up the vast majority of computational costs to produce the results in this paper.
@@ A.3 Bigram Subnetwork Selection
Thus in our analyses in §4.2, §5, and §6, for GPT-2 (re-trained), GPT-2 small, and Pythia 160M, we use the bigram subnetwork trained with $\lambda=10$. For GPT-2 large and Pythia 1B, we use the bigram subnetwork trained with $\lambda=100$. The parameter counts and bigram surprisal correlations for these subnetworks are reported in Table 1. As noted in §6, for optimal subnetworks, we consider the sparsest optimal subnetwork that still contains more active parameters than the bigram subnetwork selected for that model. In general, we find that our results remain qualitatively similar for different $\lambda$, as long as the bigram subnetworks are reasonably sparse (e.g. $10\%$ of parameters active).
@@
A.4 Importance of Trained Token Embeddings
However, keeping only the token embedding and unembedding parameters (and layernorm parameters) does not result in high bigram surprisal correlations (Table 2; “embeddings, empty subnetwork”). In other words, the embeddings do not recreate the $W_{U}W_{E}$ embedding-unembedding bigram structure described for zero-layer Transformers in Elhage et al. (2021). This is not entirely surprising, but it indicates that the bigram subnetwork parameters we find in Transformer layers are still necessary for high bigram correlations. For example, these non-embedding subnetwork parameters might be responsible for triggering the bigram information encoded in the embedding and unembedding parameters, and thus these non-embedding parameters are still important components of the bigram subnetwork.
@@ A.6 Checkpoint Results for Other Models
[…]
@@ A.9 Optimal Subnetwork Results for Other Models
In Table 3, we report the actual and expected (by chance) parameter overlaps between the bigram subnetwork and the optimal subnetwork for each model, as described in §6.2 (with bigram subnetworks selected as in §A.3). Concretely, when considering the parameter overlap between two subnetworks $C_{0}$ and $C_{1}$, we generate 10K random subnetwork pairs (pairs of parameter masks) that randomly select the same number of parameters as $C_{0}$ and $C_{1}$ respectively within each parameter block. Then, we count the number of random pairs whose parameter overlap is greater than or equal to the parameter overlap of $C_{0}$ and $C_{1}$. This gives us the probability that two random subnetworks would have equal or greater overlap than $C_{0}$ and $C_{1}$, after accounting for the number of parameters kept per block in $C_{0}$ and $C_{1}$. As in Pythia 1B in the main text (§6.2), for all models, the actual overlaps between bigram subnetworks and optimal subnetworks are far larger than would be expected by chance. Specifically, in all cases, none of the 10K randomly generated pairs have greater parameter overlap than the bigram subnetwork and the optimal subnetwork (and thus $p<0.0001$).
@@
[…]